OpenStudy (anonymous):

The total number of fungal spores can be found using an infinite geometric series where a1 = 10 and the common ratio is 3. Find the sum of this infinite series that will be the upper limit of the fungal spores. 495 990 1,485 The series is divergent

OpenStudy (anonymous):

The surface area of an oil spill gets 105% larger every day, represented by the function s(x) = (1.05)x − 1. On the first day, it covered an area of 31 square meters. Which function would be used to find the area of the oil spill on the 15th day? s(x) = (1.05)31 − 1 s(x) = (1.05)x − 31 s(x) = 1.05(31)x − 1 s(x) = 31(1.05)x − 1 @ganeshie8